OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? A. 576 B. 240 C. 144 D. 960

OpenStudy (gschibby):

384/0.4 = 960 ;)

OpenStudy (***[isuru]***):

if the amount of people who visit that site is"x" the the amount who download that special son is 60x/100 = 0.6x so the total amount of people = 0.6x + people who didn't download that song hence 60% download that song the remaining amount is 40% the equation in the problem says... x = 0.6x + 384 so... the amount who didn't download that song is 384 and 384 = 40% of the total amount... so 40x/100 = 384 x = (384*100)/40 = 960 which mean the total amount who visit that site in that month is 960 and the total amount who download that song is... 960 x (60/100 ) = 576 hope this will help ya!!